OpenStudy (anonymous):

*EASY QUESTION I KNOW BUT IM TO LAZY !NEED HELP*Sarah is selling lemonade on the street corner. Every day around 24 to 46 cars pass by the street corner. If 1 in 4 cars that pass stop to purchase lemonade, and each person that stops spends between $1 and $4, about how much will Sarah make per day?

OpenStudy (anonymous):

lol take the extreme two cases

OpenStudy (anonymous):

?

OpenStudy (valpey):

Depends on the cost of lemonade but if you assume "around 24 to 46 cars" means the average is 35 cars per day and you assume "spends between $1 and $4" means the average spent is $2.50 then: 35 cars * 25% * $2.50 =

OpenStudy (anonymous):

WHAT ARE YOU ANSWER CHOICES

OpenStudy (anonymous):

$3 $47 $54 $21

OpenStudy (valpey):

35*0.25*2.5 = ?

OpenStudy (anonymous):

OK YOU ITS NOT 3 SO YOU CAN ELIMATE THAT SO IT IS EITHER $21, $47, $54, I THINK EITHER $21 OR $47

OpenStudy (anonymous):

IM GEUSSING 21

OpenStudy (anonymous):

ME TWO BECAUSE I GOT 18.97 SO IF YOU KEEP ROUNDING IT WOULD BE $21 NO HIGH OR NO LES CUZ 47 IS TO HIGH AN 3 IS TO LOW

OpenStudy (anonymous):

OK SEE YA LTER BOUT TO GO BUYEEEEEEE

OpenStudy (dumbcow):

its 21...use averages average number of cars per day is : (24+46)/2 = 35 1 in 4 buy --> 35/4 = 8.75 average spent is: (1+4)/2 = 2.50 --> 8.75 * 2.50 = 21.875